given the function g(v)=3v+4, find g(5), g(-7), and g(6.7).
find the slope of (5,3) and (8,-6)
what is the domain of g(x)=10
---
3-7x
find the indicated outputs for f(x)=5x 2squared -5x
f(0)=
f(-1)=
f(2)=

Answer by nyc_function(2741)

BIG TIP: Post one question at a time or else no one will reply.
I will do: Given the function g(v) = 3v+4, find g(5), g(-7), and g(6.7).
Replace v with each given value simplify individually.
Let v = 5
g(5) = 3(5) + 4
g(5) = 15 + 4
g(5) = 19
===============
Let v = -7
g(-7) = 3(-7) + 4
g(-7) = -21 + 4
g(-7) = -17
===============
Let v = 6.7
g(6.7) = 3(6.7) + 4
g(6.7) = 20.1 + 4
g(6.7) = 24.1